OpenStudy (anonymous):

I need help, so that's why Im gonaa ask this question: You have two numbers, with a difference of 2. If you divide them however, you get 2.2(with a bar on top of the second 2). What are the two answers. Please give me like an equation, and then the answer, thank you soooo much whoever answers this. (No thank you to the people who saw this and turned their head :) )

OpenStudy (anonymous):

wah, no one's helping.. : (

OpenStudy (freckles):

let's call the numbers a and b so a-b=? and a/b=?

OpenStudy (anonymous):

a-b = 2 a/b= 2.22222222(repeating) or to make it more simple, a/b= 2 and two-ninths

OpenStudy (freckles):

right so we have: \[a-b=2 \\ \frac{a}{b}=\frac{20}{9}\] You can solve the first equation for a by saying a=2+b now replace the a in the second equation with 2+b

OpenStudy (freckles):

\[\frac{2+b}{b}=\frac{20}{9}\] solve for b

OpenStudy (anonymous):

hhmmm I must be stupid

OpenStudy (freckles):

why?

OpenStudy (anonymous):

erm... lemme try that.... can't u say that the b/b is equal to 1 and say that 1+ wait no.....